11 11 TEX L^{A}TEX

Linux User Group HOWTO

Kendall Grant Clark <mailto:kclark@cmpu.net> v.1.6.2, 24 April 1998

The Linux User Group HOWTO is a guide to founding, maintaining, and growing a Linux User Group.

Contents

1 Introduction 2 1.1 Purpose . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 1.2 Other sources of information . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

2 What is a Linux User Group? 2 2.1 What is Linux? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 2.2 How is Linux unique? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 2.3 What is a user group? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 2.4 Summary . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

3 What LUGs are there? 4 3.1 Lists of LUGs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 3.2 Solidarity versus convenience . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

4 What does a LUG do? 5 4.1 Linux advocacy . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 4.2 Linux education . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 4.3 Linux support . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

4.3.1 Users . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 4.3.2 Consultants . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 4.3.3 Businesses, non-profit organizations, and schools . . . . . . . . . . . . . . . . . . . . . 9 4.3.4 Free software development . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 4.3.5 Linux Movement . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11

4.4 Linux socializing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11

5 Local LUG activities 12 5.1 Meetings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 5.2 Online resources . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13

6 Practical suggestions 14 6.1 LUG support organizations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 6.2 Founding a LUG . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 6.3 Maintaining and growing a LUG . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15

7 Legal and political issues 16 7.1 Legal issues . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 7.2 United States . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 7.3 Canada . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 7.4 Political issues . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17

7.4.1 People have different feelings about free software. . . . . . . . . . . . . . . . . . . . . . 18 7.4.2 Nonprofit organizations and money don’t mix terribly well. . . . . . . . . . . . . . . . 18

8 About this document 19 8.1 Terms of use . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 8.2 New versions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 8.3 Please contribute to this HOWTO . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 8.4 Document history . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 8.5 Acknowledgements . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20

1 Introduction

1.1 Purpose

The Linux User Group HOWTO is intended to serve as a guide to founding, maintaining, and growing a Linux User Group.

Linux is a freely-distributable implementation of Unix for personal computers, servers and workstations. It was developed on the i386 and now supports i486, Pentium, Pentium Pro, and Pentium II processors, as well as x86-clones from AMD, Cyrix, and others. It also supports many SPARC, DEC Alpha, Power-PC/PowerMac, Motorola 68x0 Mac/Amiga machines.

1.2 Other sources of information

If you want to learn more about Linux, the Linux Documentation Project <http://sunsite.unc.edu/LDP/> is a good place to start.

For general information about computer user groups, please see the Association of PC Users Groups <http:

//www.apcug.org/>.

2 What is a Linux User Group?

2.1 What is Linux?

In order to appreciate and understand fully the significant role of LUGs in the Linux Movement, it is important to understand what makes Linux unique among computer operating systems.

Linux as an operating system is very efficient and very powerful. But, Linux as an idea about how software ought to be developed is even more powerful. Linux is a free operating system: it is licensed under the GNU Public License. The source code is freely available to anyone who wants it and always will be. It is developed by a unstructured group of programmers from around the world, under the technical direction of Linus Torvalds and other key developers. Linux is a world-wide movement without any central structure, bureaucracy, or entity to control, coordinate, or otherwise direct its affairs. While this situation is a powerful part of the appeal and technical quality of Linux as an computer operating system, it can make for inefficient allocation of human resources, ineffective and even detrimental advocacy, public relations, user education and training.

2.2 How is Linux unique?

This loose structure is not likely to change with regard to Linux as a software project. And it’s a good thing, too. Linux works precisely because people are free to come and go as they please: free programmers are happy programmers are effective programmers.

But this loose structure can make the average Linux user’s life a little complicated–especially if that user isn’t a programmer by profession or by vocation. Who does she call for support, training, or education? How does she know the kinds of uses for which Linux is well-suited?

In large part local LUGs provide the answers to these kinds of question. This is why LUGs are a crucial part of the Linux Movement. Because there is no “regional office” of the Linux Corporation in your town or village or metropolis, the local LUG takes on many of the same roles that a regional office does for a large multi-national corporation..

Linux is unique because it does not have, nor is it burdened by, a central structure or bureaucracy to allocate its resources, train its users, or provide support for its products. These jobs get done in a variety of ways: the Internet, consultants, VARs, support companies, colleges and universities. But, increasingly, in many places around the globe, they get done by a local LUG.

2.3 What is a user group?

Computer user groups, at least in the United States, are not a new phenomenon; in fact, they played an important role in the history of the personal computer. The personal computer arose in large part to satisfy the demand of electronics, Ham Radio, and other hobbyist user groups, as well as trade shows and swap meets, for affordable, personal access to computing resources. Of course eventually giants like IBM discovered that the PC was a good and profitable thing, but the impetus for the PC came from the people, by the people, and for the people.

In the United States, user groups have changed, and many for the worse, with the times. The financial woes of the largest user group ever, the Boston Computer Society <http://www.bcs.org/> have been well-reported; but all over the U.S. most of the big PC user groups have seen a decline in real membership. American user groups in their heyday concentrated on the production of newsletters, the maintenance of shareware and diskette libraries, meetings, social events, and, sometimes, even Bulletin Board Systems. With the advent of the Internet, however, many of the services that user groups once provided were transferred to things like CompuServe, AOL, and the Web.

The rise of Linux, however, coincided with and was intensified by general public’s “discovery” of the Internet. As the Internet grew more popular, so did Linux: the Internet brought new users, developers, and vendors to the Linux Movement.

So just when traditional PC user groups were declining because of the Internet’s popularity, this popularity propelled Linux forward, creating new demand for new user groups dedicated exclusively to Linux. To give just one indication of the ways in which a LUG is different than a traditional user group, I call the reader’s attention to a curious fact: traditional user groups have had to maintain a fairly tight control over the kinds of software that its users copy and trade at its meetings. While illegal copying of commercial software certainly occurred at these meetings, it was officially discouraged and for good reason.

At a LUG meeting, however, this entire mindset simply does not apply. Far from being the kind of thing that a LUG ought to discourage, the free copying of Linux itself ought to be one of the primary activities of a LUG. In fact there is anecdotal evidence that traditional user groups sometimes have a difficult time adapting to the fact that Linux can be freely copied as many times as one needs or wants.

2.4 Summary

In order for the Linux Movement to continue to flourish, the proliferation and success of local LUGs, along with other factors, is an absolute requirement. Because of the unique status of Linux, the local LUG must provide some of the same functions that a “regional office” provides for large computer corporations like IBM, Microsoft, or Sun. LUGs can and must train, support, and educate Linux users, coordinate Linux consultants, advocate Linux as a computing solution, and even serve as a liason to local media outlets like newspapers and television.

3 What LUGs are there?

Since this document is meant as a guide not only to maintaining and growing LUGs but also to founding them, it would be well before we go much further to determine what LUGs there are.

3.1 Lists of LUGs

There are several lists of LUGs available on the Web. If you want to found a local LUG, one of the first things to do is to determine where the nearest LUG is. Your best bet may be to join a LUG that is already established in your area rather than founding a new one.

As of the mid-1997, there are LUGs in all 50 states, the District of Columbia, and 26 other countries, including India, Russia, and most of Western and Eastern Europe.

Note: the biggest untapped computing market on the planet, China, does not yet appear to have a LUG, and India, the second most populous country on the planet, has only a few.

• Finding Groups of Linux Users Everywhere <http://www.ssc.com/glue/groups/>

   • LUG List Project <http://www.nllgg.nl/lugww/>
   • LUG Registry <http://www.linux.org/users/index.html>

It appears that the GLUE list is more comprehensive for American LUGs, while the LUG List Project offers more comprehensive international coverage.

3.2 Solidarity versus convenience

While the lists of LUGs on the Web are well-maintained, it is likely that they do not list every LUG. In addition to consulting these lists, I suggest, if you are considering founding a LUG, that you post a short message asking about the existence of a local LUG to comp.os.linux.announce <news:comp.os.linux. announce>, comp.os.linux.misc <news:comp.os.linux.misc>, or an appropriate regional Usenet hierarchy. If there isn’t a LUG already in your area, then posting mesages to these groups will alert potential members of your plans.

If you plan to found a local LUG, you should carefully balance convenience against solidarity. In other words, if there is a LUG in your metropolitan area, but on the other side of the city, it may be better to start a new group for the sake of convenience. But it may be better to join the pre-existing group for the sake of unity and solidarity. Greater numbers almost always means greater power, influence, and efficiency. While it might be nice to have two groups of 100 members each, there are certain advantages to one group of 200 members. Of course if you live in a small town or village, any group is better than no group at all.

The point is that starting a LUG is an arduous undertaking, and one that ought to be entered into with all the relevant facts, and with some appreciation of the effect on other groups.

4 What does a LUG do?

The goals of local LUGs are as varied as the locales in which they operate. There is no master plan for LUGs, nor is this document meant to supply one. Remember: Linux is free from bureaucracy and centralized control and so are local LUGs.

It is possible, however, to identify a core set of goals for a local LUG:

• advocacy

• education

• support

• socializing

Each local LUG will combine these and other goals in a unique way in order to satisfy the unique needs of its membership.

4.1 Linux advocacy

The urge to advocate the use of Linux is as natural to computer users as is eating or sleeping. When you find something that works and works well, the natural urge is to tell as many people about it as you can. The role of LUGs in Linux advocacy cannot be overestimated, especially since the wide-scale commercial acceptance of Linux which it so richly deserves has not yet been achieved. While it is certainly beneficial to the Linux Movement each and every time a computer journalist writes a positive review of Linux, it is also beneficial every time satisfied Linux users tell their friends, colleagues, employees or employers about Linux.

There is effective advocacy and there is ineffective carping: as Linux users, we must be constantly vigilant to advocate Linux in such a way as to reflect positively on both the product, its creators and developers, and our fellow users. The Linux Advocacy mini-HOWTO, available at the Linux Documentation Project, gives some helpful suggestions in this regard. Suffice it to say that advocacy is an important aspect of the mission of a local LUG.

There may come a time when Linux advocacy is pretty much beside the point because Linux has more or less won the day, when the phrase “No one ever got fired for using Linux” becomes a reality. Until that time, however, the local LUG plays an indispensable role in promoting the use of Linux. It does so because its advocacy is free, well-intentioned, and backed up by organizational commitment. If a person comes to know about Linux through the efforts of a local LUG, then that person, as a new Linux user, is already ahead of the game: she is already aware of the existence of an organization that will help her install, configure, and even maintain Linux on whatever computers she is willing to dedicate to it.

New Linux users who are already in contact with a local LUG are ahead of those whose interest in Linux has been piqued by a computer journalist, but who have no one to whom to turn to aid them in their quest to install, run, and learn Linux.

It is, therefore, important for local LUGs to advocate Linux because their advocacy is effective, wellsupported, and free.

4.2 Linux education

Not only is it the business of a local LUG to advocate the use of Linux, it may also turn its efforts to training its members, as well as the computing public in its area, to use Linux and associated components. In my own estimation, the goal of user education is the single most important goal a LUG may undertake. Of course, as I have already pointed out, LUGs are perfectly free to organize themselves and their activities around any of these, or other, goals. I believe, however, that LUGs can have the greatest impact on the Linux Movement by educating and training Linux users.

Local LUGs may choose to undertake the goal of education simply because there is no other local entity from which a Linux user may receive technically-oriented education. While it is certainly the case that universities, colleges, and junior colleges are increassingly turning to Linux as a way to educate their students, both efficiently and cheaply, about Unix-like operating systems, some Linux users are either unable or unwilling to register for courses in order to learn Linux. For these users the local LUG is a valuable resource for enhancement or creation of advanced computer skills: Unix-like system administration, system programming, support and creation of Internet and Intranet technologies, etc.

In an ironic twist, many local LUGs are even sharing the burden of worker training with large corporations. Every worker at Acme Corp that expands her computer skills by participating in a local LUG is one less worker Acme Corp has to train or pay to train. Even though using and administering a Linux PC at home isn’t the same as administering a corporate data warehouse, call center, or similar high-availability facility, it is light years more complex, more rewarding, and more educational than using and administering a Windows 95 PC at home. As Linux itself advances toward things like journalling filesystems, high-availability, realtime capacity, and other high-end Unix features, the already blurry line between Linux and the “real” Unixes will get even more indistinct.

Not only is such education a form of worker training, but it will also serve, as information technology becomes an increasingly vital part of the global economy, as a kind of community service. In most metropolitan areas in the United States, for example, it is possible for a local LUG to take Linux into local schools, small businesses, community and social organizations, and other non-corporate environments. This accomplishes the task of Linux advocacy and also helps train the general public about Linux as a Unix-like operating system. As more and more of these kinds of organizations seek to establish an Internet presence or provide dial-in access to their workers, students, and constituents, the opportunities arise for local LUGs to participate in the life of their community by educating it about a free and freely-available operating system. This kind of community service allows the average Linux user to emulate the kind of generosity that has characterized Linux, and the free software community, from the very beginning. Most Linux users can’t program like Linus Torvalds, but we can all all give our time and abilities to other Linux users, the Linux community, and the broader community in which work and live.

Linux is a natural fit for these kinds of organization because deploying it doesn’t commit them to expensive license, upgrade, or maintenance fees. Because Linux is also technically elegant and economical, it runs very well on the the kinds of disposable hardware that corporations typically cast off and that non-profit organizations are only too happy to use. As more and more people discover every day, that old 486 collecting dust in the closet can do real work if someone will install Linux on it.

In addition, Linux education has a cumulative effect on the other goals of a local LUG, in particular the goal of Linux support discussed below. Better Linux education means better Linux support. The more people that a LUG can count on to reach its support goals, the easier support becomes and, therefore, the more of it can be done. The more new and inexperienced users a local LUG can support and eventually educate about Linux, the larger and more effective the LUG can become. In other words, if a LUG focuses solely on Linux support to the neglect of Linux education, the natural barriers to organizational growth will be more restrictive. If only two or three percent of the members of a LUG take upon themselves the task of supporting the others, the growth of the LUG will be stifled. One thing you can count on: if new and inexperienced users don’t get the help with Linux they need from a local LUG, they won’t participate in that LUG for very long. If a larger percentage of members support the others, the LUG will be able to grow much larger. Linux education is the key to this dynamic: education turns new Linux users into experienced ones.

Free education about free Linux also highlights the degree to which Linux is part and parcel of the free software Community. So it seems appropriate that local LUGs focus not solely on Linux education but also education about all of the various software systems and technologies that run under Linux. These include, for instance, the GNU suite of programs and utilities, the Apache Web server, the XFree86 implementation of X Windows, TeX, LaTeX, etc. Fortunately the list of free software that runs under Linux is a long and diverse one.

Finally, Linux is a self-documenting operating environment; in other words, if we don’t write the documentation, nobody is going to do it for us. Toward that end, make sure that LUG members are well aware of the Linux Documentation Project <http://sunsite.unc.edu/LDP/>, which can be found at mirrors worldwide. Consider providing an LDP mirror for the local Linux community and for LUG members. Also make sure to publicize—through comp.os.linux.announce, the LDP, and other pertinent sources of Linux information— any relevant documentation that is developed by the LUG: technical presentations, tutorials, local FAQs, etc. There is a lot of Linux documentation produced in LUGs that doesn’t benefit the worldwide Linux community because no one outside the LUG knows about it. Don’t let the LUGs efforts in this regard go to waste: it is highly probable that if someone at one LUG had a question or problem with something, then people at other LUGs around the world will have the same questions and problems.

4.3 Linux support

Of course for the desperate newbie the primary role of a local LUG is Linux support. But it is a mistake to suppose that Linux support only means technical support for new Linux users. It can and should mean much more.

Local LUGs have the opportunity to support:

• users

• consultants

• businesses, non-profit organizations, and schools

• the Linux Movement

4.3.1 Users

The most frequent complaint from new Linux users, once they have gotten Linux installed, is the steep learning curve which is not at all unique to Linux but is, rather, a characteristic of all modern Unixes. With the steepness of the learning curve, however, comes the power and flexibility of a complex operating system. A local LUG is often the only resource that a new Linux user has available to help flatten out the learning curve.

But even if a new Linux user doesn’t know it yet, she needs more than just technical support: Linux and the free software worlds are both rapidly moving targets. The local LUGs form an invaluable conduit of information about Linux and other free software products. Not only does Linux lack a central bureaucracy, but it also for the most part lacks the kind of journalistic infrastructure from which users of other computer systems benefit. The Linux Movement does have resources like Linux Journal <http://www.ssc.com/lj/> and Linux Gazette <http://www.ssc.com/lg/>, but many new Linux users are unaware of these resources. In addition, as monthly publications they are often already out of date about bugfixes, security problems, patches, new kernels, etc. This is where the local LUG as a source and conduit of timely information is so vital to new and experienced Linux users alike.

For example, until a new Linux user knows that the newest kernels are available from ftp.kernel.org <ftp: //ftp.kernel.org> or that the Linux Documentation Project usually has newer versions of Linux HOWTOs than a CD-based Linux distribution, it is up to the local LUG, as the primary support entity, to be a conduit of timely and useful information.

In fact it may be just a bit misleading to focus on the support role that local LUGs provide to new users: intermediate and advanced users also benefit from the proliferation of timely and useful tips, facts, and secrets about Linux. Because of the complexity of Linux, even advanced users often learn new tricks or techniques simply by becoming involved in a local LUG. Sometimes they learn about software packages they didn’t know existed, sometimes they just remember that arcane vi command sequence they’ve not used since college.

4.3.2 Consultants

It is, I think, rather obvious to claim that local LUGs ought to be in the business of supporting new Linux users. After all, if they’re not supposed to be doing that, what are they to do? It may not be as obvious that local LUGs can play an important role in supporting local Linux consultants. Whether they do Linux consulting full-time or only part-time, consultants can be an important part of a local LUG. How can the LUG support them?

The answer to that question is just the answer to another question: what is it that Linux consultants want and need? They need someone for whom to consult. A local LUG provides the best way for those who offer Linux consulting to find those who need Linux consulting. The local LUG can informally broker connections between consulting suppliers and consulting consumers simply by getting all, or as many as possible, of the people interested in Linux in a local area together and talking with one another. How LUGs do that will occupy us below. What is important here is to point out that LUGs can and should play this role as well. The Linux Consultants HOWTO is an important document in this regard, but it is surely the case that only a fraction of the full-time and part-time Linux consultants worldwide are registered in the Consultants HOWTO.

The relationship is mutually beneficial. Consultants aid LUGs by providing experienced leadership, both technically and organizationally, while LUGs aid consultants by putting them in contact with the kinds of people who need their services. New and inexperienced users gain benefit from both LUGs and consultants since their routine or simple requests for support are handled by LUGs gratis, and their complex needs and problems—the kind that obviously require the services of a paid consultant—can be handled by the consultants whom the local LUG helps them contact.

The line between support requests that need a consultant and those that do not is sometimes indistinct; but in most cases the difference is clear. While a local LUG doesn’t want to gain the reputation for pawning new users off unnecessarily on consultants–as this is simply rude and very anti-Linux behavior–there is no reason for LUGs not to help broker contacts between the users who need consulting services and the professionals who offer them.

Please see Martin Michlmayr’s Linux Consultants HOWTO <http://sunsite.unc.edu/LDP/HOWTO/ Consultants-HOWTO.html> for an international list of Linux consultants.

4.3.3 Businesses, non-profit organizations, and schools

LUGs also have the opportunity to support local businesses and organizations. This support has two aspects. First, LUGs can support businesses and organizations that want to use Linux as a part of their computing and IT efforts. Second, LUGs can support local businesses and organizations that develop for Linux, cater to Linux users, support or install Linux, etc.

The kinds of support that LUGs can provide to local businesses that want to use Linux as a part of their computing operations isn’t really all that different from the kinds of support LUGs give to individuals who want to run Linux at home. For example, compiling the Linux kernel doesn’t really vary from home to business. Supporting businesses using Linux, however, may mean that a LUG needs to concentrate on commercial software that runs on Linux, rather than concentrating solely on free software. If Linux is going to continue to maintain its momentum as a viable computing alternative, then it’s going to take software vendors who are willing to write for and port to Linux as a commercially-viable platform. If local LUGs can play a role in helping business users evaluate commercial Linux solutions, then more software vendors will be encouraged to consider Linux in their development and planning.

This leads us directly to the second kind of support that a local LUG can give to local businesses. Local LUGs can serve as a clearing house for the kind of information that is available in very few other places. For example:

• Which local ISP is Linux-friendly?

• Are there any local hardware vendors that build Linux PCs?

• Does anyone sell Linux CDs locally?

Maintaining and making this kind of information public not only helps the members of a local LUG, but it also helps Linux-friendly local businesses as well, and it encourages them to continue to be Linux-friendly. It may even, in some cases, help contribute to a competitive atmosphere in which other businesses are encouraged to become Linux-friendly too.

4.3.4 Free software development

Finally, LUGs may also support the Linux Movement by soliciting and organizing charitable giving. Chris Browne <mailto:cbbrowne@hex.het> has thought about this issue as much as anyone I know, and he contributes the following.

Chris Browne on free software philanthropy A further involvement can be to encourage sponsorship of various Linux-related organizations in a financial way. With the multiple millions <http://counter.li.org> of Linux users, it would be entirely plausible for grateful users to individually contribute a little. Given millions of users, and the not unreasonable sum of a hundred dollars of “gratefulness” per Linux user ($100 being roughly the sum not spent this year upgrading a Microsoft OS), that could add up to hundreds of millions of dollars towards development of improved tools and applications for Linux.

A users group can encourage members to contribute to various “development projects.” If it has some form of “charitable tax exemption” status, that can encourage members to contribute directly to the group, getting tax deductions as appropriate, with contributions flowing on to other organizations.

It is appropriate, in any case, to encourage LUG members to direct contributions to organizations with projects and goals that they individually wish to support.

This section lists possible candidates. None are explicitly being recommended here, but the list can represent useful “food for thought.” Many are registered as charities in the United States, thus making U.S. contributions tax deductible.

Here are organizations with activities particularly directed towards development of software that works with Linux:

   • Linux International Project Sponsorship Fund <http://www.li.org/About/Fund/Welcome.html>
   • Debian/Software In the Public Interest <http://www.debian.org/donations.html>
   • Free Software Foundation <http://www.fsf.org/help/donate.html>
   • The XFree86 Project <http://www.xfree86.org/donations.html>

Contributions to these organizations has the direct effect of supporting the creation of freely redistributable software usable with Linux. Dollar for dollar, such contributions almost certainly have greater effect on the Linux community as a whole than any other specific kind of spending.

There are also organizations that are less directly associated with Linux that may nonetheless be worthy of assistance, such as:

• League for Programming Freedom <http://www.lpf.org> This is not a Linux-specific organization; they are involved in general advocacy activities that touch on people involved with software development. Involvement in this organization represents something closer to involvement in a “political lobby” group. There is somewhat of a “USA bias;” there are nonetheless international implications, and the international community as often follows the American lead in computing-related matters as vice-versa.

• The LaTeX3 Project Fund The TeX Users Group (TUG) <http://www.tug.org> is working on the “next generation” version of the LaTeX publishing system, known as LaTeX3. Linux is one of the platforms on which TeX and LaTeX are best supported. Donations for the project can be sent to:

          TeX Users Group
          P.O. Box 1239
          Three Rivers, CA 93271-1239
          USA

or, for those in Europe,

          UK TUG
          1 Eymore Close
          Selly Oaks
          Burmingham B29 4LB
          UK
   • Project Gutenberg <http://www.promo.net/pg/lists/list.html>

Their purpose is to make freely available in electronic form the texts of out-of-copyright books. This isn’t directly a “Linux thing,” but it seems fairly worthy, and they actively encourage platform independence, which means that their “products” are quite usable with Linux.

4.3.5 Linux Movement

I have referred throughout this HOWTO to something I call the Linux Movement. There really is no better way to describe the international Linux phenomenon than to call it a movement: it isn’t a bureaucracy, but it is organized; it isn’t a corporation, but it is important to businesses all over the world. The best way for a local LUG to support the international Linux movement is to work to insure that the local Linux community is robust, vibrant, and growing. Linux is developed internationally, which is easy enough to see by reading /usr/src/linux/MAINTAINERS. But Linux is also used internationally. And this ever-expanding user base is the key to Linux’s continued success. And that is where the local LUG plays an incalculably important role.

The strength of the Linux Movement internationally is the simple fact that Linux offers unprecedented computing power and sophistication for its cost and for its freedom. The keys are value and independence from proprietary control. Every time a new person, group, business, or organization has the opportunity to be exposed to Linux’s inherent value the Linux Movement grows in strength and numbers. Local LUGs can make that happen.

4.4 Linux socializing

The last goal of a local LUG that I will mention here is socializing. In some ways this is the most difficult goal to discuss because it is not clear how many or to what degree LUGs engage in it. While it would be strange to have a local LUG that didn’t engage in the other goals, there very well may be local LUGs somewhere in the world for which socialization isn’t an important consideration.

It seems, however, that whenever two or three Linux users get together fun, highjinks, and, often, beer are sure to follow. Linus Tovalds has always had one enduring goal for Linux: to have more fun. For hackers, kernel developers, and Linux users, there’s nothing quite like downloading a new kernel, recompiling an old one, twittering with a window manager, or hacking some code. It is the sheer fun of Linux that keeps many LUGs together, and it is this kind of fun that leads many LUGs naturally to socializing.

By “socializing” here I mean primarily sharing experiences, forming friendships, and mutually-shared admiration and respect. There is another meaning, however, one that social scientists call socialization. In any movement, institution, or human community, there is the need for some process or pattern of events in and by which, to put it in Linux terms, newbies are turned into hackers. In other words, socialization turns you from “one of them” to “one of us”.

For armed forces in the U.S. and in most countries, this process is called boot camp or basic training. This is the process whereby civilians are transformed into soldiers. The Linux movement has analogous requirements. It is important that new Linux users come to learn what it means to be a Linux user, what is expected of them as a member of an international community, the special vocabulary of the Linux movement, its unique requirements and opportunities. This may be as simple as how Linux users in a partcicular locale pronounce “Linux”. It may be as profound as the ways in which Linux users should advocate, and the ways in which they should, more importantly, refrain from advocating Linux. Linux socialization, unlike ‘real world’ socialization, can occur on mailing lists and Usenet, although the efficacy of the latter is constantly challenged precisely by poorly socialized users. In my view, socialization and socializing are both done best in the company of real, flesh-and-blood fellow human beings, and not by incorporeal voices on a mailing list or Usenet group.

5 Local LUG activities

In the previous section I focused exclusively on what LUGs do and what they ought to be doing. In this section the focus shifts to practical strategies for accomplishing these goals. There are, despite the endless permutations of form, two basic things that local LUGs do: first, they meet together in physical space; second, they communicate with each other in cyberspace. Everything or nearly everything that LUGs do can be seen in terms of meetings and online resources.

5.1 Meetings

As I said above, physical meetings are synonymous with LUGs and with most computer user groups. LUGs have these kinds of meetings:

• social

• technical presentations

• informal discussion groups

• user group business

• Linux installation

• configuration and bug-squashing What do LUGs do at these meetings?

• Install Linux for newbies and strangers

• Teach members about Linux

• Compare Linux to other operating systems

• Teach members about the software that runs on Linux

• Discuss the ways in which Linux can be advocated

• Discuss the importance of the Free Software Movement

• Discuss the business of the user group

• Eat, drink, and be merry

5.2 Online resources

The commercial rise of the Internet coincided roughly with the rise of Linux, and the latter in large part owes something to the former. The Internet has always been an important asset for Linux development. It is no different for LUGs. Most LUGs have web pages if not whole Web sites. In fact, I am not sure how else to find a local LUG but to check the Web. It makes sense, then, for a local LUG to make use of whatever Internet technologies they can appropriate: Web sites, mailing lists, gopher, FTP, e-mail, WAIS, finger, news, etc. As the world of commerce is discovering, the Internet can be an effective way to advertise, inform, educate, and even sell. The other reason that LUGs make extensive use of Internet technologies is that it is the very essence of Linux to provide a stable and rich platform for the deployment of these technologies. So not only do LUGs benefit from, say, the establishment of a Web site because it advertizes their existence and helps organize their members, but in deploying these technologies, the members of the LUG are provided an opportunity to learn about this technology and see Linux at work. Some LUGs that use the Internet effectively:

• Atlanta Linux Enthusiasts <http://www.ale.org/>

• North Texas Linux Users Group <http://www.ntlug.org/>

• Boston Linux and Unix <http://www.blu.org/>
• Colorado Linux Users and Enthusiasts <http://spot.elfwerks.com/~clue/>
• BLUG - BHZ Linux Users Group (Brazil) <http://www.bhz.ampr.org/~linux/>

• Ottawa Carleton Linux Users Group <http://www.oclug.on.ca/>

• Provence Linux Users Group <http://www.pipo.com/plug/>

• Duesseldorf Linux Users Group <http://www.hsp.de/~dlug/>

• Linux User Group Austria <http://www.luga.or.at/>
• Israeli Linux Users Group <http://www.linux.org.il/>
• Tokyo Linux Users Group <http://www.twics.co.jp/~tlug/>
• Linux in Mexico <http://www.linux.org.mx/>

• Netherlands Linux Users Group (NLLGG) <http://www.nllgg.nl/>

• St. Petersburg Linux User Group <http://ethereal.ru/~mbravo/spblug/index.html>

• Linux User Group of Singapore <http://www.lugs.org.sg/>

• Victoria Linux User Group <http://www.linux.victoria.bc.ca/>
• Essex Linux User Group <http://www.epos.demon.co.uk/>
• Turkish Linux User Group <http://www.linux.org.tr/>

• Linux User Group of Rochester <http://www.lugor.org/>

• Korean Linux Users Group <http://www.linux-kr.org>

Please let me know if your LUG uses the Internet in an important or interesting way; I’d like this list to include your group.

6 Practical suggestions

Finally, I want to make some very practical, even mundane, suggestions for anyone wanting to found, maintain, or grow a LUG.

6.1 LUG support organizations

There are several organizations that offer assistance to local LUGs.

GLUE Groups of Linux Users Everywhere is a user group coordination and support program started by SSC, the same people who publish Linux Journal. The GLUE program <http://www.ssc.com/glue/> is an inexpensive way for a local LUG to provide some benefits to its membership.

Linux Systems Labs LSL <http://www.lsl.com/> offers their Tri-Linux Disk set (Three Linux distributions on four CDs: Red Hat, Slackware, and Debian) to LUGs for resale at a considerable discount.

Linux Mall User Group Program Sponsored by WorkGroup Solutions, the Linux Mall User Group Program <http://www.LinuxMall. com/usergrp.program.html> offers a range of benefits for participating User Groups. LUGs are also free to participate in Linux Mall’s Referral Program <http://www.LinuxMall.com/mallrfr.html> as well.

Cleveland Linux User’s Group Owns the Internet domain, lug.net. They will provide your LUG an Internet domain name at lug.net: your-LUG-name-or-citylug.net. More information may be found at LUG.NET <http://www.lug. net/> or by e-mailing Jeff Garvas.

Red Hat Software’s User Group Program Assists LUGs to develop and grow. More information may be found at Red Hat Web site <http:

//www.redhat.com/redhat/rhug.html>

6.2 Founding a LUG

• Determine the nearest pre-existing LUG

• Announce your intentions on comp.os.linux.announce and on an appropriate regional hierarchy

• Announce your intention wherever computer users are in your area: bookstores, swap meets, cybercafes, colleges and universities, corporations, Internet service providers, etc.

• Find Linux-friendly businesses or institutions in your area that may be willing to help you form the LUG

• Form a mailing list or some means of communication between the people who express an interest in forming a LUG

• Ask key people specifically for help in spreading the word about your intention to form a LUG

• Solicit space on a Web server to put a few HTML pages together about the group

• Begin looking for a meeting place

• Schedule an initial meeting

• Discuss at the initial meeting the goals for the LUG

6.3 Maintaining and growing a LUG

• Make the barriers to LUG membership as low as possible

• Make the LUG’s Web site a priority: keep all information current, make it easy to find details about meetings (who, what, and where), and make contact information and feedback mechanisms prominent

• Install Linux for anyone who wants it

• Post flyers, messages, or handbills wherever computer users are in your area

• Secure dedicated leadership

• Follow Linus’s benevolent dictator model of leadership

• Take the big decisions to the members for a vote

• Start a mailing list devoted to technical support and ask the “gurus” to participate on it

• Schedule a mixture of advanced and basic, formal and informal, presentations

• Support the software development efforts of your members

• Find way to raise money without dues: for instance, selling Linux merchandise to your members and to others

• Consider securing formal legal standing for the group, such as incorporation or tax-exempt status • Find out if your meeting place is restricting growth of the LUG

• Meet in conjunction with swap meets, computer shows, or other community events where computer users—i.e., potential Linux converts—are likely to gather

• Elect formal leadership for the LUG as soon as is practical: some helpful officers might include President, Treasurer, Secretary, Meeting Host (general announcements, speaker introductions, opening and closing remarks, etc.), Publicity Coordinator (handles Usenet and e-mail postings, local publicity), and Program Coordinator (organizes and schedules speakers at LUG meetings)

• Provide ways for members and others to give feedback about the direction, goals, and strategies of the LUG • Support Linux and Free Software development efforts by donating Web space, a mailing list, or FTP

site

• Establish an FTP site for relevant software

• Archive everything the LUG does for the Web site

• Solicit “door prizes” from Linux vendors, VARs, etc. to give away at meetings

• Give credit where credit is due

• Join SSC’s GLUE (Groups of Linux Users Everywhere) but be aware they charge a membership fee • Submit your LUG’s information to all of the Lists of LUGs

• Publicize your meetings on appropriate Usenet groups and in local computer publications and newspapers

• Compose promotional materials, like Postscript files, for instance, that members can use to help publicize the LUG at workplaces, bookstores, computer stores, etc.

• Make sure you know what LUG members want the LUG to do

• Release press releases to local media outlets about any unusual LUG events like an Installation Fest, Net Day, etc.

• Use LUG resources and members to help local non-profit organizations and schools with their Information Technology needs

• Advocate the use of Linux zealously but responsibly

• Play to the strengths of LUG members

• Maintain good relations with Linux vendors, VARs, developers, etc.

• Identify and contact Linux consultants in your area

• Network with the leaders of other LUGs in your area, state, region, or country to share experiences, tricks, and resources

• Keep LUG members advised on the state of Linux software—new kernels, bugs, fixes, patches, security advisories—and the state of the Linux world at large—new ports, trademark and licensing issues, where Linus is living and working, etc.

• Notify the Linux Documentation Project—and other pertinent sources of Linux information—about the documentation that the LUG produces: technical presentations, tutorials, local HOWTOs, etc.

7 Legal and political issues

7.1 Legal issues

7.2 United States

There is a strong case to be made for formal organization of local LUGs. I will not make that case here. If, however, you are interested in formally organizing your local LUG, then this section will introduce you to some of the relevant issues.

Note: this section should not be construed as competent legal counsel. These issues require the expertise of competent legal counsel; you should, before acting on any of the statements made in this section, consult an attorney.

There are at least two different legal statuses that a local LUG in the United States may attain:

1. incorporation as a non-profit entity

2. tax-exemption

Although the relevant statutes differ from state to state, most states allow user groups to incorporate as non-profit entitites. The benefits of incorporation for a local LUG may include limitations of liability of LUG members and volunteers, as well as limitation or even exemption from state corporate franchise taxes.

While you should consult competent legal counsel before incorporating your LUG as a non-profit entity, you can probably reduce your legal fees if you are acquainted with the relevant issues before consulting with an attorney. I recommend the Non-Lawyers Non-Profit Corporation Kit (ISBN 0-937434-35-3).

As for the second status, tax-exemption, this is not a legal status so much as a judgment by the Internal Revenue Service. It is important for you to know that incorporation as a non-profit entity does not insure that the IRS will rule that your LUG is to be tax-exempt. It is possible to have a non-profit corporation that is not also tax-exempt.

The IRS has a relatively simple document that explains the criteria and process for tax-exemption. It is Publication 557: Tax-Exempt Status for Your Organization. It is available as an Adobe Acrobat file from the IRS’s Web site. I strongly recommend that you read this document before filing for incorporation as a non-profit entity. While becoming a non-profit corporation cannot insure that your LUG will be declared tax-exempt by the IRS, there are ways to incorporate that will prevent the IRS from declaring your LUG to be tax-exempt. Tax-Exempt Status for Your Organization clearly sets out the necessary conditions for your LUG to be declared tax-exempt.

Finally, there are resources available on the Internet for non-profit and tax-exempt organizations. Some of the material is probably relevant to your local LUG.

7.3 Canada

Thanks to Chris Browne for the following comments about the Canadian situation.

The Canadian tax environment strongly parallels the US environment, in that the “charitable organization” status confers similar tax advantages for donors over mere “not for profit” status, while requiring that similar sorts of added paperwork be filed by the “charity” with the tax authorities in order to attain and maintain certified charity status.

7.4 Political issues

Chris Browne <mailto:cbbrowne@hex.net> has the following to say about the kinds of inter-LUG political dynamics that often crop up.

7.4.1 People have different feelings about free software.

Linux users are a diverse bunch. As soon as you try to put a lot of them together, there are some problem issues that can come up.There are those that are nearly political radicals that believe that all software, always, should be “free.” Because Caldera charges quite a lot of money for their distribution, and doesn’t give all profits over to (pick favorite advocacy organization), they must be “evil.” Ditto for Red Hat or S.u.S.E. Keep in mind that all three of these companies have made and continue to make significant contributions to free software.

Others may figure that they can find some way to highly exploit the “freeness” of the Linux platform for their fun and profit. Be aware that many users of the BSD UNIX variants consider that their licenses that do permit companies to build “privatized” custom versions of their OSes are preferable to the “enforced permanent freeness” of the GPL as applied to Linux. Do not presume that all people promoting this sort of view are necessarily greedy leeches.

If these people are put together in one place, disagreements can occur.

Leaders should be clear on the following facts:

• There are a lot of opinions about the GPL and how it is supposed to work. It is easy to misunderstand both the GPL and alternative licensing schemes.

• Linux benefits from contributions from many places, and can support some freeloaders, particularly if this encourages more people to get involved, thus pulling in further contributors.

• Many significant contributions have been made to Linux by commercial enterprises. Examining the sources to the Linux kernel, and notable subsystems such as XFree86 and GCC show a surprising number of commercial contributors.

• Commercial does not always imply “better,” but it also does not always imply “horrible.”

The main principle can be extended well beyond this; computer “holy wars” have long been waged over the virtues of one system over another, whether that be (in modern day) between Linux, other UNIX variants, and Microsoft OSes, or between the “IBM PC” and the various Motorola 68000-based systems, or between the many 8 bit systems of the 1970s. Or of KDE versus GNOME.

A wise LUG leader will seek to smooth over such differences, rather than inciting them. LUG leaders must have thick skins.

There will be disagreements at some point as diverse views collide with one another, and leaders must be able to cope with this, resolving disagreements rather than contributing to the problem.

7.4.2 Nonprofit organizations and money don’t mix terribly well.

It is important to be quite careful in dealing with finances in a nonprofit organization of any sort. In businesses, where profitable flows of monies are the goal, people are not typically too worried about “nagging details” such as possible misspending of immaterial sums of money.

The same cannot be said about nonprofit organizations. Some people are involved for reasons of principle, and can easily give minor problems inordinate attention. And the potential for wide participation at business meetings correspondingly expands the potential for inordinate attention to be drawn to things.

As a result, it is probably preferable for there to not be a membership fee for a LUG, as that provides a specific thing for which people can reasonably demand accountability. Fees that are not collected cannot, by virtue of the fact that they don’t exist, be misused.

If there is a lot of money and/or other such resources floating around, it is important for the user group to be accountable to its members for it.

In a vital, growing group, there should be more than one person involved. In troubled nonprofit organizations, financial information is often tightly held by someone who will not willingly relinquish control of funds. Ideally, there should be some rotation of duties in a LUG including that of control of the finances.

Regular useful financial reports should be made available to those that wish them. A LUG that maintains an official “charitable status” for tax purposes will have to file at least annual financial reports with the local tax authorities, which would represent a minimum financial disclosure for the purposes of the members.

With the growth of Linux-based financial software, it should be readily possible to create reports on a regular basis. With the growth of the Internet, it should even be possible to publish these on the World Wide Web.

8 About this document

8.1 Terms of use

Copyright (c) 1997 by Kendall Grant Clark. This document may be distributed under the terms set forth in the LDP license at http://sunsite.unc.edu/LDP/COPYRIGHT.html <http://sunsite.unc.edu/

LDP/COPYRIGHT.html>.

8.2 New versions

New versions of the Linux User Group HOWTO will be periodically uploaded to various Linux WWW and FTP sites, principally my homepage <http://www.ntlug.org/~kclark/> and the Linux Documentation

Project <http://sunsite.unc.edu/LDP/>

8.3 Please contribute to this HOWTO

I welcome questions about and feedback on this document. Please send them to me at kclark@cmpu.net. I am especially interested in hearing from leaders of LUGs from around the world. I would like to include real-life examples of the things described here. I would also like to include a section on LUGs outside the United States, since this HOWTO as it stands now is rather US-centric. Please let me know if your group does things that should be mentioned in this HOWTO.

8.4 Document history

• 1.0 released on 13 July 1997

• 1.1: expanded online resources section

• 1.3: added LUG Support Organizations and expanded the Legal and Organizational Issues section • 1.3.1: general editing for clarity and conciseness

• 1.4: general editing, added new LUG resources

• 1.4.1: general editing for clarity

• 1.5: added some resources, some discussion of LUG documentation, also general editing

• 1.5.1: changed Web location for this document and author’s email address.

• 1.5.2: new copyright and license

• 1.5.3: miscellaneous edits and minor re-organizations

• 1.6: added Chris Browne’s material: Linux philanthropic donations and LUG political considerations • 1.6.1: very minor additions

• 1.6.2: minor corrections

8.5 Acknowledgements

I want to thank all the great people I’ve met and worked with during the time I’ve served as President of the North Texas Linux Users Group. They helped inspire me to use Linux full-time. The best thing about Linux really is the people you meet. I especially want to thank Chris Browne <mailto:cbbrowne@hex.net> for describing the situation with non-profit and charitable groups in Canada, his thoughts on financial donations as a way to participate in Linux and the free software movement, and his ideas about the kinds of political issues that may arise within LUGs. In addition, the following people have made helpful comments and suggestions:

• Hugo van der Kooij

• Greg Hankins

• Charles Lindahl

• Rick Moen

• Jeff Garvas

• James Hertzler